# DF40: Toward Next-Generation Deepfake Detection

### v1

@@ A.3.2 Full Experimental Results and Discussion
A.3.3 Further Analysis Results
Deeper Analysis of the Influence of Data Domains and Forgery Methods
We analyze logits and confidence distributions for real and fake classes using Xception and SBI models, visualized in Fig. 6. (1) For Xception: (i) Same Domain and Method: Real and fake class distributions have small variance and high mean, with minimal ambiguity in confidence. (ii) Cross Method, Same Domain: Fake class distribution has increased variance and lower mean, while real class remains relatively unchanged. (iii) Cross Domain, Same Method: Real class distribution has increased variance and lower mean, while fake class remains relatively unchanged, causing a slight performance decrease. (iv) Cross Method and Domain: Both real and fake class distributions have significantly increased variance and lower means, decreasing overall performance. This verifies Finding-6. (2) For SBI: It shows a different behavior than Xception. The real class distribution consistently has a small variance and high mean, while the fake class distribution has a larger variance and lower mean. This suggests that SBI learns a robust distribution for real classes, generalizing well in the FF domain using the FS method (0.81 AUC). However, when the real class distribution changes ($e.g.$, in the CDF domain), performance drops significantly (to 0.679 AUC). This indicates that SBI might operate like an "anomaly detection method" based on deviations from the learned real class distribution.
Training Config CLIP Other Detectors
Image Size 224, 224 256, 256
Weight Initialization CLIP Pre-trained ImageNet Pre-trained
Optimizer Adam Adam
Base Learning Rate 1e-5 2e-4
Weight Decay 5e-4 5e-4
Optimizer Momentum B1, B2=0.9, 0.999 B1, B2=0.9, 0.999
Batch Size 32 32
Training Epochs 15 15
Learning Rate Schedule None, Constant None, Constant
Flip Probability 0.5 0.5
Rotate Probability 0.5 0.5
Rotate Limit [-10, 10] [-10, 10]
Blur Probability 0.5 0.5
Blur Limit [3, 7] [3, 7]
Brightness Probability 0.5 0.5
Brightness Limit [-0.1, 0.1] [-0.1, 0.1]
Contrast Limit [-0.1, 0.1] [-0.1, 0.1]
Quality Lower 40 40
Quality Upper 100 100
@@ A.3.3 Further Analysis Results
A.4 Dataset Publication
A.4.1 Links
We provide a private direct link to the dataset is shared with the reviewers at https://anonymous.4open.science/r/DF40_6E0E/ to keep the single-blind review, where we give the OneDrive download link, associated codes, and visual examples for illustrating partial forgery images/videos of DF40. We will provide the official website, download links, and license for the dataset after publication. Our dataset with documentation and associated codes (for reproduction of our experimental results) is maintained under the following license: https://creativecommons.org/licenses/by-nc/4.0/.
A.4.2 Hosting Platform & Access to Dataset
In the review process, we host our dataset on the OneDrive account, where reviewers and AC can access the fake videos and images of our dataset. After publication, we will host our dataset on an official website with detailed instructions and official download links. Also, after publication, a dataset request form (Google form) to monitor and restrict the free use of our deepfake dataset. In this manner, everyone has to fill out the dataset request form, which we will manually screen to limit misuse.
A.4.3 Licensing
Our DF40 dataset is available under the Creative Commons NonCommercial license (CC BY-NC). More details can be seen in https://creativecommons.org/licenses/by-nc/4.0/).

### v2

@@ A.4.2 Full Experimental Results and Discussion
A.5 Further Analysis Results
A.5.1 More Results of Different Fake Regions
We create a self-reconstruction image (SRI) for each original real image, where the SRI images act as fakes with a global forgery pattern (entire image synthesis). We then perform a blending operation to replace the reconstructed face with the original face, creating a localized fake (L-SRI), similar to other typical localized face forgery methods. We consider five common types of fake regions: whole, face, mouth, nose, and eyes. Results in Tab. 17 quantitatively verify the significance of the fake region in detection.
Table 17: Ablation study regarding the impact of the fake region.
  Whole Face Mouth Nose Eye
0.952 / 0.953 0.730 / 0.700 0.477 / 0.454 0.483 / 0.458 0.470 / 0.454
Table 18: Results of evaluating three models ($i.e.$, I3D, CLIP-base, and CLIP-large) trained on different variants of the proposed DF40 on other previous deepfake datasets.
  Model Training Set Testing Set
FR (FF) 0.684 0.917 0.846 0.374 0.895 0.876 0.807
FS (FF) + FR (FF) 0.764 0.970 0.926 0.994 0.928 0.953 0.929
CLIP-base FS (FF) 0.823 0.965 0.678 0.985 0.585 0.932 0.945
FR (FF) 0.649 0.669 0.572 0.457 0.682 0.625 0.947
EFS (FF) 0.553 0.618 0.604 0.634 0.539 0.653 0.590
DF40 (FF) 0.799 0.932 0.698 0.973 0.683 0.889 0.937
CLIP-large FS (FF) 0.905 0.989 0.777 0.991 0.608 0.963 0.965
FR (FF) 0.698 0.756 0.614 0.487 0.738 0.626 0.983
EFS (FF) 0.616 0.733 0.634 0.730 0.571 0.772 0.739
DF40 (FF) 0.891 0.978 0.753 0.973 0.717 0.927 0.986
  Model Training Set Testing Set
ADM BigGAN GLide MidJourney SD-v4 SD-v5 Vqdm Wukong Avg.
FR (FF) 0.695 0.595 0.230 0.505 0.437 0.417 0.549 0.519 0.493
EFS (FF) 0.912 0.992 0.906 0.591 0.729 0.72 0.956 0.713 0.815
DF40 (FF) 0.911 0.967 0.736 0.571 0.630 0.614 0.882 0.660 0.746
A.5.2 Generalization to Non-face Domain AIGCs
A.5.3 Generalization to External/Previous Deepfake Datasets
We also evaluate the model trained on DF40 to detect previous deepfake datasets. Results in Tab. 18 show that (1) For the I3D detection method, training on FS (FF) and FR (FF) combined (FS (FF) + FR (FF)) results in better performance across most testing sets compared to training on FS (FF) or FR (FF) individually. This suggests that combining face-swapping and face-reenactment datasets for training improves the generalization capabilities of the I3D method. (2) Among the CLIP-base detection methods, training on DF40 (FF) leads to better performance across most testing sets compared to training on FS (FF), FR (FF), or EFS (FF) individually. This indicates that incorporating diverse deepfake techniques in the training set improves the generalization capabilities of the CLIP-base method. (3) The same trend is observed for the CLIP-large detection methods, where training on DF40 (FF) results in better performance across most testing sets compared to training on FS (FF), FR (FF), or EFS (FF) individually. This further supports the importance of training on diverse deepfake techniques to enhance the generalization capabilities of the detection methods. (4) The CLIP-large detection method generally outperforms the CLIP-base method across most testing sets when trained on the same training sets. This suggests that the larger model size of CLIP-large contributes to better deepfake detection performance.
Figure 15: Logits and confidence analysis for comparing CLIP-base and CLIP-large. We show that CLIP-large can learn more robust real-people features, thereby generalizing well in unseen forgeries.
A.5.4 Comparative Analysis of CLIP-base and CLIP-large for Deepfake Detection
t-SNE visualization:
Logits Distribution Analysis:
Figure 16: t-SNE visualizations for comparing CLIP-base and CLIP-large. We show that CLIP-large can learn more robust real-people features, thereby generalizing well in unseen forgeries.
Figure 17: t-SNE visualization and logits analysis for analyzing the model, which is trained on FR (FF) and tested on StyleCLIP, achieving only about 0.2 AUC. See the text for our analysis and more details.
Figure 18: Frequency visualizations (std) of both the training and testing phases, for real and fake, respectively.
A.5.5 Analysis of "Anomalous Values" in Experiments
In our previous experimental results summarized in Tab. 5 (protocol-3), we observed an anomalous outcome where the Xception model, trained on FF (FR), achieved an AUC of less than 0.2 on the StyleCLIP dataset. Given that an AUC of 0.5 indicates performance by chance, this result can be particularly perplexing. Here, we would like to delve deeper and clarify this observation in detail, as follows:
Result explanation:
A very low prediction logits of the fake class make the 0.006 AUC possible: Conceptually, AUC can be formulated as $AUC=P(P_{p}>P_{n})$, where $P_{p}$ and $P_{n}$ represent the probabilities of positive (fake, labeled as 1) and negative (real, labeled as 0) samples, respectively. In our case, the AUC of 0.006 can be attributed to the model assigning very low logits (e.g., 10-7) to the fake class compared to the real class (e.g., 10-2), as evidenced by the logit distribution in Fig. 17. According to the AUC formulation, when $P_{p}$ (logits for fake samples) is predominantly lower than $P_{n}$ (logits for real samples), the AUC approaches 0. Intuitively, this result indicates that the model perceives the fake test samples (in this case, StyleCLIP) as more likely to be real than the actual real samples. As a result, the ROC curve lies close to the x-axis, leading to an AUC of 0.006.
Underlying reason:
Unaligned resolution between the real and fake classes introduces a model bias: To understand the underlying reason behind it, we provide further analysis from the resolution perspective. Specifically, we observe that the training real data (FF++) obviously contains more high-frequency components (higher resolution) compared to the training fake data (FR). This noticeable resolution gap between real and fake data could be inevitably captured by the model [56], leading the model to potentially adopt a trivial solution: "Low resolution implies fake." In contrast, during testing, the fake data (StyleCLIP) contains significantly more high-frequency components (higher resolution) than the testing real data used for testing. We provide a frequency visualization as evidence to validate this claim. As shown in Fig. 18, the testing fake (StyleCLIP) contains much more high-frequency details than the testing real. This bias leads the model to make an inconsistent decision, resulting in the anomaly in AUC. In our case, intuitively, since the biased model "observes" StyleCLIP carrying more high-frequency details with higher resolution than the real, the model then "thinks" StyleCLIP is more real than the actual real, eventually resulting in the 0.006 AUC.
Figure 19: Confidence distribution comparison between Xception (image model) and I3D (video model). We show that I3D has less uncertainty in making real/fake classification decisions, while Xception has higher uncertainty.
A.5.6 Comparison of Video Model and Image Model
Here, we would like to provide a deeper analysis comparing a video method (I3D) and an image method (Xception) using confidence distribution analysis.
Specifically, we obtain the prediction logits for each testing image from both models and apply the softmax function to obtain the confidence scores (ranging from 0 to 1). The visualization results of the confidence distribution for Xception and I3D can be seen in Fig. 19. As illustrated, we observe that I3D exhibits less uncertainty when making decisions for detection, while Xception shows higher uncertainty. This may suggest that there is limited discriminative information for Xception to perform detection based on a single frame. In contrast, the temporal information (many frames within a video) provides more informative clues for I3D’s detection. This could be attributed to the fact that most (advanced) FS and FR methods cannot ensure perfect consistency at the temporal level, but they can maintain high realism at the image level, making detection more challenging for image-based detectors.
Figure 20: t-SNE visualization for illustrating the detection results of CLIP-large on the MidJourney dataset. We show that (fake) people of different races and genders can group into different clusters, highlighting that CLIP-large has already learned rich information about the real people’s attributions.
A.5.7 Visualizations of the Real People Features.
Here, we explore which types of people tend to cluster in the MJ dataset and show the reasons behind this clustering. In Fig. 7 of our main paper (the result of CLIP large on MidJourney), we indeed notice that different clusters of fake images tend to form distinct groups rather than grouping together. This is an interesting observation, and we provide further exploration and analysis of this phenomenon:
First, we create fake generated data for MidJourney using different demographic compositions. We select people from different races and genders, including four races (white, black, east Asian, and south Asian) for both males and females, resulting in eight types in total. Second, we label all eight types with different labels. Along with the real label (real data is randomly selected from FFHQ), we have nine labels in total. We use t-SNE for visualization. Results in Fig. 20 show that the clusters in t-SNE correspond to different attributes/demographics. We observe that people with different races and genders are implicitly separated by the CLIP-large model. This indicates that CLIP-large contains rich real-world information from its large-scale pre-training, making CLIP-large able to have accurate representations of races and genders. In contrast, Results in Fig. 7 (the result of Xception on MidJourney) shows that Xception cannot separate people of different races or genders. This suggests that Xception might only learn forgery-specific low-level signals without the ability to learn semantic facial attributes.
A.5.8 Artifacts of Deepfake Forgeries in Frequency
Brief Introduction:
Mean Analysis:
We observe that many GAN-generated images, such as those from StyleGAN, often exhibit a checkerboard or grid-like pattern of bright spots on the mean heatmap. This phenomenon occurs because of the upsampling methods used in the GAN architecture, which can introduce such artifacts during the image generation process. These patterns are typically a result of the transposed convolution operations, which can cause uneven overlaps and thus create these visible grid artifacts. Moreover, it appears that different forgery methods can share similar patterns. For instance, face-swapping techniques like SimSwap also exhibit these GAN-like checkerboard artifacts. This similarity indicates that these methods generate the entire image using AI without a blending process. This contrasts with other Deepfake methods, such as those used in Celeb-DF, which involve blending and thus do not produce the same checkerboard artifacts.
Std Analysis:
@@ A.5.8 Artifacts of Deepfake Forgeries in Frequency
A.6 Dataset Publication
A.6.1 Hosting Platform & Links
We host our DF40 dataset on GitHub (https://github.com/YZY-stack/DF40). Our project page is https://yzy-stack.github.io/homepage_for_df40/. All the codes, datasets, and checkpoints are released at the above links. Our dataset with documentation and associated codes (for reproduction of our experimental results) is maintained under the Creative Commons NonCommercial license (CC BY-NC). More details can be seen in https://creativecommons.org/licenses/by-nc/4.0/).
A.6.2 Controlled Access:
To control the random distribution, we have already implemented a controlled access system (a Google form) for the DF40 dataset, where users may require authorization or meet specific prerequisites to gain access (see https://docs.google.com/forms/d/e/1FAIpQLSedrbobvH3JtCZR7wDzH4S9olikmwL67x7AawckgRXKF_T2Mg/viewform). This approach, similar to protocols used with previous deepfake datasets like FF++ and Celeb-DF, can help restrict indiscriminate access. Additionally, acknowledging the evolving nature of threats, one of our strategies is the continuous refinement and adaptation of the tools and detection methods associated with the DF40 dataset. A static dataset is more vulnerable to exploitation; therefore, keeping our platform adaptive is essential.
A.6.3 Discrimination, Bias, and Fairness:
Please note that the DF40 fake data is created entirely based on the real data, meaning that the demographic composition of the datasets is the same as the original datasets. We chose this setting to align with previous works. Specifically, most prior deepfake detection studies train their models on four deepfake methods from FF++ and test them on one fake method from CDF. This consideration motivated us to choose FF++ and CDF as the real data sources for generating corresponding fake data for alignment.
However, as shown in [74], these datasets might not be fair enough to create a completely unbiased detector without any explicit constraints and supervision. To address this potential concern, we have proposed several solutions to address the fairness issue, including: a. We have added the facial attribute annotations (such as race and gender) for each video on our GitHub page (github.com/YZY-stack/DF40/tree/main/Annotations_for_Facial_Attributes). By providing access to these attributes, users can develop fairer and less biased deepfake detectors using our datasets; b. Most existing studies on fairness are based on the FF++ and CDF datasets, the same datasets we utilized for developing ours. This suggests that existing fairness methods can be also used in our dataset. In summary, the existing fairness studies such as [31] that enforce fairness constraints based on the annotations of each video in FF++, can directly apply their methods to our dataset for improved fairness.